# Expanding Deep Learning-based Sensing Systems with Multi-Source Knowledge Transfer

### v1

@@ 5.2.2. Vaired Target Architectures
HaKT 79.27 2517 59.30 1279 45.12 6.95 63.96 525
@@
Table 3. Detailed performance comparison on the HARBox dataset in the multi-round expansion setting.
@@
HaKT 77.74 296 81.05 534 79.42 741 78.85 946 79.27 2517
5.2.3. Varied Portion of Labeled Data
The portion of labeled data is varied to evaluate the robustness of HaKT. As shown in Figure 7, HaKT outperforms or achieves comparable performance to the baselines in 13 out of 16 cases. Specifically, on the HARBox dataset, when $\gamma$ is 0.10, 0.15, and 0.20, HaKT achieves accuracy improvements of 5.04%, 3.77%, and 4.56%, respectively. This is achieved by leveraging both labeled and unlabeled data to select high-quality models and generate accurate pseudo-labels for unlabeled data through sample-wise knowledge fusion. In few cases, the best baseline slightly outperforms HaKT (e.g., when $\gamma=0.05$ on HARBox). This may be due to the use of a fixed threshold $b$ in the adaptive learners, which may not yield an optimal weight $\alpha$ across different values of $\gamma$. To address this, we plan to explore dynamic thresholding to further enhance HaKT’s robustness. Besides, we observe that the performance of the baselines is generally lower than reported in their original papers. This discrepancy is likely due to the more challenging experiment settings, which involve limited labels and large heterogeneities that HaKT is specifically designed to address.
During the one-time expansion, HaKT reduces the average communication traffic by 39%, 36%, 8.3%, and 37.5% compared with baselines that use all source domain models on the HARBox, ImageNet-R, NinaPro, and AD datasets, respectively. Additionally, the source model execution cost in HaKT is reduced by approximately 40% when $\eta=60\%$, demonstrating the efficiency of HaKT in the expansion process. Besides, when the number of selected models $N_{p}$ varies from one to three on the HARBox dataset, the performance achieved by HaKT increases from 75.43% to 80.57%, which is 4.9% and 5.0% higher than the best baselines, respectively. This improvement highlights the effectiveness of HaKT in aggregating knowledge from multiple models.
5.3. Result Comparison on Multi-Round Sensing Systems Expansion
We further present the effectiveness of HaKT in scaling sensing systems in a multi-round expansion setting.
5.3.1. Data Split and Training Details
The domains in each dataset are randomly divided into five groups, denoted as $\{G_{i},i=0,\cdots,4\}$. Detailed information about the groups is provided in Table 4. In round $j$, the domains in $\{G_{i},i=0,\cdots,j-1\}$ serve as the source domains, while the domains in $G_{j}$ are the target domains, with their model skeleton randomly selected from the model libraries. Once the models in $G_{j}$ are ready for use, they are incorporated as source domains in the subsequent round $j+1$, sharing their knowledge with new targets for further expansion. For instance, during round 2 expansion on the HARBox dataset, the source domains include 60 users from $G_{0}$ and $G_{1}$, whose knowledge is used to build models for 20 users in $G_{2}$. After round 2 expansion, the system scales from 60 to 80 users, and the models in $G_{2}$ are subsequently leveraged to construct models for $G_{3}$ along with $G_{0}$ and $G_{1}$. The parameter $\gamma$ is set to 10%. All other settings are kept consistent with Section 5.2.
Table 4. Group information for the four datasets.
HARBox ImageNet-R NinaPro AD
Groups [40, 20, 20, 20, 20] [4, 1, 1, 1, 1] [6, 1, 1, 1, 1] [8, 2, 2, 2, 2]
5.3.2. Multi-Round System Expansions
Table 2 presents the performance in different rounds of expansion. The average accuracy achieved by HaKT is 3.9%, 1.6%, 0.2%, and 5.8% higher compared with the best baselines on the four datasets. Due to the space limit, we present the detailed results of the four rounds on the HARBox dataset in Table 3, and similar results have been observed for other datasets. In Table 3, HaKT outperforms the best baselines by 4.2%, 3.2%, 3.1%, and 4.3% in accuracy. In round $j$, HaKT provides better customized models for group $G_{j}$ since it selects better source models, fuses knowledge in finer granularity, and injects knowledge dynamically according to their quality. In contrast, DistillWeighted assigns fixed weights to models, which may not fully capture the varying quality of each source on different samples. Moreover, as models trained in earlier rounds can be used later, models trained with HaKT consistently transfer higher-quality knowledge to the target models in subsequent rounds, leading to better performance on the new targets.
In Table 2 and Table 3, as the system scales, the increase in the number of source domains results in larger traffic. HaKT effectively reduces this communication burden across all datasets. For instance, on the HARBox dataset, HaKT reduces the communication traffic by 1.5 GB compared with DistillWeighted. Note that the communication traffic of LEAP and MEHLSoup is not directly comparable to HaKT, as these methods are constrained to using a limited subset of source domains with architectures that align with the target domains. This limitation stems from their inability to handle model heterogeneity, which also contributes to their inferior performance. Overall, the results demonstrate HaKT’s superior effectiveness and efficiency in scaling sensing systems.
5.4. Ablation Study
5.4.1. Effectiveness of the Designs.
Table 5. Ablation study in the one-time expansion setting. SwKF and AKI refer to the Sample-wise Knowledge Fusion and Adaptive Knowledge Injection.
Design HARBox ImageNet-R NinaPro AD
HaKT 80.57 57.87 49.38 67.33
w/o FbCS 73.86 56.86 47.75 63.83
w/o CAJS 79.92 56.45 44.95 67.17
w/o SwKF 69.05 56.14 41.75 60.17
w/o AKI 66.73 53.29 40.66 48.50
Direct Training 70.31 51.97 45.50 60.33
Table 5 shows that the Feature-based Coarse Selection (FbCS) and Centroids-Accuracy Joint Selection can enhance performance by up to 6.7% and 4.4% in accuracy, respectively. The reason is that the statistical features and high-level representations based on labeled and unlabeled data could accurately reflect the domain similarity and the source model effectiveness. Besides, we observe that a larger improvement from the Coarse Selection typically leads to a smaller incremental gain from the Fine-grained Selection. This may occur because the lightweight features of specific modalities are sufficient for selecting high-quality models. However, the combination of both selection methods demonstrates stronger generalizability across different tasks and modalities. The Sample-wise Knowledge Fusion achieves an 11.5% accuracy improvement on HARBox, which is attributed to the sample-wise weights learned by the attention-based mixer could more effectively combine predictions from the source models. Additionally, the Adaptive Knowledge Injection improves accuracy by at least 2.1%, highlighting the benefit of selectively storing fused predictions and adjusting the weight of the distillation loss based on their quality. The performance gains of HaKT over the Direct Training approach demonstrate the advantages of leveraging source domain models for training target models.
Optimizing the Model Selection Overhead. The impact of the FbCS on communication overhead is illustrated in Figure 8(a) by varying its selection ratio, $\eta$. Figure 8(a) shows that as $\eta$ decreases from 1.0 (without FbCS) to 0.03 (without Centroids-Accuracy Joint Selection), the communication traffic consistently decreases because fewer models are selected for transmission. The model execution cost also decreases as fewer source models are chosen for encoding data in the target domain. Additionally, Figure 8(a) shows that the accuracy achieved by the target model increases and then slightly decreases as $\eta$ decreases. This pattern occurs because, when $\eta$ is large, the FbCS filters out less useful models. However, when $\eta$ becomes too small, the coarse selection inadvertently discard some high-quality models.
Figure 8. (a) The impact of the selection ratio in the Feature based Coarse Selection. (b) The impact of the Cost-effective Adaptation and Low-cost Joint Training.
Table 6. Performance comparison of different model selection methods on the HARBox dataset.
Random Accuracy PARC HaKT
Acc (%) 56.21 58.99 64.59 65.70
Traffic (MB) 13.0 858 858 526
Table 7. Performance comparison of different knowledge fusion methods on the HARBox dataset.
Metrics Nearest Equal AccDistill Weighted HaKT
[…]
Acc (%) 43.48 46.81 49.19 45.63 65.70
    The accuracy of the pseudo labels.
Optimizing the Training Overhead. Figure 8(b) presents the training overhead on the ImageNet-R dataset. The high memory usage is due to the large size of vision models. The Cost-effective Adaptation, which partially tune the selected models during training, lead to a 2.0$\times$ reduction in memory usage and a 2.3$\times$ reduction in training time due to fewer parameters being optimized. Similarly, incorporating Low-cost Joint Training reduces the per-epoch training time from 64.1s to 37.8s. Overall, HaKT achieves significant reductions in both training time (4.6$\times$) and memory usage (2.7$\times$), indicating a more energy-efficient training process.
5.4.2. Alternatives in Model Selection
Several alternative model selection methods are compared with the selection approach of HaKT in Table 6. The knowledge transfer process in HaKT is applied to all selection methods. Accuracy and the PARC criteria achieve better performance compared with random selection. However, both methods rely on the labeled data, making them less effective when presented with label scarcity. In contrast, the Efficient Model Selection Protocol in HaKT leverages both labeled and unlabeled data for selection and avoids full model transmission, resulting in a 1.1% accuracy improvement while using only 61.3% traffic of the communication expense.
5.4.3. Alternatives in Knowledge Fusion
The Sample-wise Knowledge Fusion method is compared with other knowledge fusion methods in Table 7. Nearest represents no fusion, where one single model is selected and used. Weighted indicates the use of the fusion method from DistillWeighted. Equal refers to assigning equal weights to all selected models. The accuracy of the pseudo labels generated by HaKT is 11.6% higher than the best alternative method. Consequently, by learning from the higher-quality fused knowledge, the target models achieve a 16.5% improvement in accuracy.
6. Discussions
Applicability to Resource-Constrained Devices. The diversity of device types in various sensing systems presents challenges for customized model training. To minimize system overhead during expansion, HaKT optimizes communication traffic through an efficient model selection protocol and reduces training memory and time with a low-cost joint training scheme, making the expansion process more feasible for edge servers. However, the complete model training process may still exceed the capabilities of battery-powered IoT devices and wearables. In such cases, offloading model training to nearby trusted edge servers or leveraging edge-cloud collaboration can serve as effective solutions (Samikwa et al., 2023; Wang et al., 2024).
Privacy Concerns during System Expansion. Most domain adaptation methods require simultaneous access to both source and target domain data, which limits their applicability in privacy-sensitive scenarios (He et al., 2023; Qu et al., 2024). In contrast, HaKT better preserves data privacy by exchanging only high-level features and models between domains. While sharing features may still carry some risk of sensitive information leakage, it is generally necessary to identify relevant domains (Bolya et al., 2021). In future work, we aim to enhance privacy further by selectively sharing non-sensitive features through methods that identify and exclude sensitive content (Qu et al., 2024).
7. Related Works
Transfer Learning. Transfer learning explores methods to apply existing models to new targets, overcoming data or task heterogeneities (Pan and Yang, 2009; Tan et al., 2018). To address data heterogeneity, domain adaptation techniques have been widely studied to align the feature distributions between source and target domain (Zhu et al., 2020; Wilson et al., 2021; He et al., 2023). However, most methods require access to both source and target domain data during training, which is not feasible in the scenario considered. Additionally, some multi-source transfer learning approaches focus on selecting source models with better generalizability (Tong et al., 2021; Agostinelli et al., 2022; Lee et al., 2019). For example, (Agostinelli et al., 2022) proposes constructing empirical predictors for model selection, which estimate the performance models could achieve after transfer. However, these methods often overlook device heterogeneity, which might prevent the direct adoption of source models in target domains.
Knowledge Distillation. In knowledge distillation, an efficient student model is trained using the knowledge of one or more teacher models, such as their predicted pseudo labels or intermediate features (Liu et al., 2020; Romero et al., 2014; Hinton et al., 2015; Vemulapalli et al., 2024). Specifically, multi-teacher distillation approaches (Borup et al., 2023; Liu et al., 2020; Zhang et al., 2022) aggregate the knowledge of multiple teachers by assigning weights, aiming to provide the student model with more accurate and comprehensive knowledge. Most knowledge distillation studies focus on a closed-set problem, where teachers with high-quality knowledge are predetermined and available to use(Hinton et al., 2015; Liu et al., 2020; Zhang et al., 2022; Borup et al., 2023). However, in the sensing system expansion problem, the knowledge from source domain models may not be directly applicable to the target domain due to data heterogeneity, leading to suboptimal performance.
Model Customization. Model customization has been extensively studied to meet specific computational and performance requirements (Wen et al., 2023; Cai et al., 2020; Ren et al., 2021). Directly training customized models on target data is less effective due to label scarcity (Ouyang et al., 2023; Xu et al., 2023). To address this, self-supervised learning methods that leverage unlabeled data have been proposed to enhance performance (Xu et al., 2021; Ouyang et al., 2022). These methods are not included as baselines for performance comparison as they are orthogonal to HaKT. Some works explore pre-deployment or post-deployment model generation techniques based on the specific requirements and conditions of target environments (Cai et al., 2020; Wen et al., 2023; Liu et al., 2018). While these methods focus on optimal architecture search in terms of latency and accuracy, HaKT emphasizes the knowledge transfer process from the selected source models to any target models that satisfy the customized needs of target domains.
8. Conclusion
To efficiently expand sensing systems, a general knowledge transfer framework, HaKT, is designed to address label scarcity and data and device heterogeneities. HaKT employs an Efficient Model Selection Protocol to identify high-quality source domain models at a low cost. The knowledge from the selected models is aggregated using Sample-wise Knowledge Fusion, which assigns different weights to each sample. The fused knowledge is then distilled into customized target models through Adaptive Knowledge Injection. Extensive experiments across various tasks, modalities, and settings demonstrate the effectiveness and efficiency of HaKT compared with state-of-the-art baselines.

### v2

@@ 5.1 Experiment Setting
HaT 79.27 2517 59.30 1279 45.12 6.95 63.96 525
[…]
5.2 Result Comparison in MRSE
We present the effectiveness of HaT in the MRSE setting. The domains in the datasets are split into four groups to simulate the learning systems expansion process in section 3.1. The details of the data split are in Appendix A.4.
Figure 5: Performance comparison with different target models.
Table 1 presents the performance across different rounds of expansion, with more detailed results provided in Appendix C. Compared to baselines that either leverage limited knowledge from source domains (LEAD and MEHLSoup) or transfer knowledge in a static manner (AccDistill, DistillNearest, and DistillWeighted), HaT delivers more effective customized models for target domains. This improvement is achieved by fusing knowledge from multiple high-quality models at a finer granularity and dynamically injecting knowledge during training. Furthermore, since models trained in earlier rounds are incorporated into subsequent rounds, HaT enables a progressive transfer of higher-quality knowledge, resulting in consistently improved performance for new target domains.
In Table 1, as the system scales, the increase in the number of source domains results in larger traffic. HaT effectively reduces this communication burden across all datasets. Note that the communication traffic of LEAP and MEHLSoup is not directly comparable to HaT, as these methods are constrained to using a limited subset of source domains with architectures that align with the target domains. This limitation stems from their inability to handle model heterogeneity, which also contributes to their inferior performance. Besides, HaT reduces the model inference cost during selection by 40%, since only a portion of source models are transmitted and executed on the target domains. Overall, the results demonstrate HaT’s effectiveness and efficiency in expanding learning systems.
Figure 6: Performance comparison with varied $\gamma$.
5.3 Result Comparisons in OTSE
We further present the performance of HaT on HARBox and ImageNet-R in the OTSE setting. The preprocessing details are in Appendix A.5. More results on the other two datasets can be found in Appendix D
Vaired Target Architectures. Figure 5 presents the results when different architectures are selected as the target models. Unlike domain adaptation and model merging methods that are limited to source models matching the target architecture (this makes them inapplicable when the required architecture, such as CPC-S for HARBox, is not available among the source models), HaT effectively utilize knowledge from source models with diverse architectures, thus achieving superior accuracy for most architectures. These results demonstrate HaT’s versatility in delivering high-quality customized models across diverse tasks and modalities.
Varied Portion of Labeled Data. As shown in Figure 6, HaT outperforms or achieves comparable performance to the baselines in most cases when the portion $\gamma$ is varied. This is achieved by adaptively fusing and injecting high-quality knowledge from the selected source models. In few cases, the best baseline slightly outperforms HaT. This may be due to the use of a fixed threshold $b$ in the adaptive learners, which may not yield an optimal weight $\alpha$ across different $\gamma$. We plan to explore dynamic thresholding to further enhance HaT’s robustness. Besides, we observe that the performance of the baselines is generally lower than reported in their original papers, which is likely due to the more challenging settings involving labels scarcity and large heterogeneities that HaT addresses.
Varied Number of Selected Models. Figure 7(a) shows that HaT consistently achieve better performance when $N_{p}$ varies, which highlights the effectiveness of HaT in fusing knowledge from multiple models. As $N_{p}$ increases, the accuracy first increases, which justifies the usage of multiple source models in HaT. When $N_{p}$ continues to increase, the accuracy drops. It might be due to the low quality knowledge in the additional models, which shows the importance of source models selection.
5.4 Ablation Study
Table 2: Ablation study in OTSE. SwKF and AKI refer to Sample-wise Knowledge Fusion and Adaptive Knowledge Injection.
Design HARBox ImageNet-R NinaPro AD
HaT 80.57 57.87 49.38 67.33
w/o FbCS 73.86 56.86 47.75 63.83
w/o CAJS 79.92 56.45 44.95 67.17
w/o SwKF 69.05 56.14 41.75 60.17
w/o AKI 66.73 53.29 40.66 48.50
Design Effectiveness. As shown in Table 2, the Feature-based Coarse Selection and Centroids-Accuracy Joint Selection enhance performance by leveraging the statistical features and high-level representations that accurately reflect the domain similarity and the source models effectiveness. The combination of both selection stages demonstrates stronger generalizability across datasets. The Sample-wise Knowledge Fusion achieves an 11.5% accuracy improvement on HARBox, which is attributed to the sample-wise weights learned by the attention-based mixer could more effectively combine predictions from the source models. Additionally, the accuracy improvement of Adaptive Knowledge Injection highlights the benefit of selectively storing fused predictions and adjusting the weight of the distillation loss based on their quality.
Figure 7: (a) The impact of the number of selected models. (b) The impact of the Cost-effective Adaptation (CeA) and Low-cost Joint Training (LcJT).
Table 3: Performance comparison of different model selection methods on the HARBox dataset.
Random Accuracy PARC HaT
Acc (%) 56.21 58.99 64.59 65.70
Traffic (MB) 13.0 858 858 526
Table 4: Performance comparison of different knowledge fusion methods on the HARBox dataset.
Metrics Nearest Equal Weighted HaT
[…]
Acc (%) 43.48 46.81 45.63 65.70
    The accuracy of the pseudo labels.
Optimizing the Training Overhead. Figure 7(b) presents the training overhead on the ImageNet-R dataset. The Cost-effective Adaptation, which partially tune the selected models during training, lead to a 2.0$\times$ reduction in memory usage and a 2.3$\times$ reduction in training time due to fewer parameters being optimized. Similarly, incorporating Low-cost Joint Training reduces the per-epoch training time from 64.1s to 37.8s. Overall, HaT achieves significant reductions in both training time (4.6$\times$) and memory usage (2.7$\times$), indicating a more energy-efficient training process.
Alternatives in Model Selection. Several alternative model selection methods are compared with the selection approach of HaT in Table 3. The knowledge transfer process in HaT is applied to all selection methods. Accuracy and the PARC criteria achieve better performance compared with random selection. However, both methods rely on the labeled data, making them less effective when presented with label scarcity. In contrast, the Efficient Model Selection Protocol in HaT leverages both labeled and unlabeled data for selection and avoids full model transmission, resulting in a 1.1% accuracy improvement while using only 61.3% traffic of the communication expense.
Alternatives in Knowledge Fusion. Different knowledge fusion methods are compared in Table 4. Nearest represents only one model is selected and used. Weighted indicates the use of the fusion method from DistillWeighted. Equal refers to assigning equal weights to all selected models. The accuracy of the pseudo labels generated by HaT is 11.6% higher than the best alternative method. Consequently, by learning from the higher-quality fused knowledge, the target models achieve a 16.5% improvement in accuracy.
Appendix A Details of Experiment Settings
A.1 Datasets.
HaT is evaluated on the four datasets in Table 5, which are further introduced as follows:
Alzheimer’s Disease (AD) (Ouyang et al., 2023). This dataset consists of Alzheimer’s Disease-related activity data collected from 16 home environments using multiple modalities. It includes 11 activity classes, such as writing and sleeping.
While HaT is evaluated on these three diverse applications, it has the potential to extend to other learning systems, such as traffic management or smart agriculture (Jiang et al., 2023; Zhu et al., 2018), which we plan to explore in future work.
Table 5: Datasets for evaluating the effectiveness and generalizability of HaT.
Dataset Class Domain (#) Task Modality Model Library
Audio, Radar
A.2 Model Libraries.
We include six different models for each dataset. For IMU data, the TPN-(S, M, L) (Saeed et al., 2019) and CPC-(S, M, L) (Haresamudram et al., 2021) models are used, with feature channels of 12, 16, 32 for TPN and 8, 12, 16 for CPC, respectively. For image processing, the model library consists of GoogleNet (Szegedy et al., 2015), MobileNet-v3 (S, L) (Howard et al., 2019), and ResNet-(18, 34, 50) (He et al., 2016). For EMG data, the ConvNet-(S, M, L) (Côté-Allard et al., 2019) models and a RNN-(S, M, L) are utilized, with feature channels of 4, 8, 12 for ConvNet and 32, 48, 56 for RNN, respectively. To handle the multi-modal data in AD, we adapt the model in (Ouyang et al., 2023) by varying the number of layers and feature dimensions, creating 5-layer ADNet-(S, M, L) and 3-layer TinyADNet-(S, M, L) models with 64, 128, 256 for ADNet and 32, 64, 96 feature channels for TinyADNet.
A.3 Baselines
HaT is compared with five most relevant baselines covering the area of knowledge distillation, domain adapation, and model aggregation:
MEHLSoup (Li et al., 2024a). MEHLSoup merges multiple source domain models with a learned mixing coefficient, which is optimized by a block coordinate gradient descent algorithm on the target domain data.
AccDistill. We select and ensemble source domain models with top-k accuracy and then transfer the knowledge from the ensembled model to the target models, leveraging the distilltion methods in (Borup et al., 2023).
Other knowledge distillation, domain adaptation, or model merging methods are not included, as they have already been outperformed by the considered baselines (Borup et al., 2023; Qu et al., 2024; Li et al., 2024a). Since LEAD and MEHLSoup, as well as other adaptation and merging methods, are unable to handle model heterogeneity, they are not directly comparable to HaT. To make both methods executable, we select source domains with architectures that match the target models as candidates. Federated learning methods are not included for comparison due to the difference in the considered scenario (See Section 3.1). Self-supervised learning methods (Xu et al., 2021; Ouyang et al., 2022) are not included as baselines as they are orthogonal to HaT and can be combined to further enhance performance.
A.4 Data Splits and Training Details in MRSE
The domains in each dataset are randomly divided into five groups, denoted as $\{G(i),i=0,\cdots,4\}$. Detailed information about the groups is provided in Table 6. In round $j$, the domains in $\{G(i),i=0,\cdots,j-1\}$ serve as the source domains $\mathcal{D}^{S}(j)$, while the domains in $G(j)$ are the target domains $\mathcal{D}^{T}(j)$. Once the models in $G(j)$ are ready for use, they are incorporated as source domains in the subsequent round $j+1$, sharing their knowledge with new targets for further expansion. For instance, during round 2 expansion on the HARBox dataset, the source domains include 60 users from $G(0)$ and $G(1)$, whose knowledge is used to build models for 20 users in $G(2)$. After round 2 expansion, the system scales from 60 to 80 users, and the models in $G(2)$ are subsequently leveraged to construct models for $G(3)$ along with $G(0)$ and $G(1)$. The model skeletons for all domains are randomly selected from TPN-(S, M, L), ResNet-(18, 34, 50), ConvNet-(S, M, L), and ADNet-(S, M, L) for the four datasets, respectively.
The source domain models are trained using supervised learning on the labeled data of each domain. For the target domains, 60% of the data is randomly selected as the training set, 20% as the validation set, and the remaining 20% as the test set. The parameter $\gamma$ is set to 10%.
Table 6: Group information for the four datasets.
HARBox ImageNet-R NinaPro AD
Groups [40, 20, 20, 20, 20] [4, 1, 1, 1, 1] [6, 1, 1, 1, 1] [8, 2, 2, 2, 2]
Figure 8: Sensitivity Analysis. (a) The impact of the selection ratio $\eta$. (b) The impact of the threshold $b$.
A.5 Data Splits and Training Details in OTSE
We also compare the results of HaT against the baselines in the OTSE setting, where one domain is randomly selected as the target domain, and the remaining domains serve as source domains. The source domain architectures are randomly selected from TPN-(S, M, L), ResNet-(18, 34, 50), ConvNet-(S, M, L), and ADNet-(S, M, L). The other settings are kept aligned with those in Section A.4. For ImageNet-R, each of the eight styles is tested separately. For the other three datasets, ten different splits are randomly generated, and the average accuracy and communication overhead are reported.
Appendix B Sensitivity Analysis
We further conduct a sensitivity analysis on two key parameters in HaT, the selection ratio $\eta$ and the threshold $b$.
Varied Selection Ratio. The impact of the Feature-based Coarse Selection (FbCS) on communication overhead is illustrated in Figure 8(a) by varying its selection ratio, $\eta$. Figure 8(a) shows that as $\eta$ decreases from 1.0 (without FbCS) to 0.03 (without Centroids-Accuracy Joint Selection), the communication traffic consistently decreases because fewer models are selected for transmission. The model execution cost also decreases as fewer source models are chosen for encoding data in the target domain. Additionally, Figure 8(a) shows that the accuracy achieved by the target model increases and then slightly decreases as $\eta$ decreases. This pattern occurs because, when $\eta$ is large, the FbCS filters out less useful models. However, when $\eta$ becomes too small, the coarse selection inadvertently discard some high-quality models.
Varied Threshold. Figure 8(b) illustrates that as the threshold $b$ increases, the average accuracy of the target models initially improves but eventually declines. A small threshold results in frequent model updates early in training, during which the aggregated predictions from the mixer are of low quality, leading to suboptimal performance. Conversely, an excessively large threshold causes the adaptive knowledge injection process to degrade into direct training with limited labeled data, thereby failing to leverage the knowledge from source models. Since the quality of selected models varies across domains and tasks, we recommend setting the threshold slightly higher than the highest accuracy of the selected models on the labeled target data. This recommendation is based on the insight that knowledge from different models can complement one another to improve overall performance.
@@ Appendix B Sensitivity Analysis
Table 7: Detailed performance comparison on the HARBox dataset in the MRSE setting.
@@
HaT 77.74 296 81.05 534 79.42 741 78.85 946 79.27 2517
(a) Varied target models.
(b) Varied portion $\gamma$.
Figure 9: Results Comparsion on NinaPro and AD in the OTSE setting. (a) The target model skeleton are varied. (b) The portion of labeled data in the target domain are varied.
Appendix C Detailed Results Comparison in the MRSE setting
Table 7 presents detailed results on each rounds of expansion on HARBox in the MRSE setting. Similar results are observed on the other datasets. HaT outperforms the best baselines by 4.2%, 3.2%, 3.1%, and 4.3% in accuracy from round 1 to 4, respectively. Besides, the traffic during expansion is significant less compared with the methods that does not have constraints in the source model architectures (LEAD and MEHLSoup only leverage source models that share the same architectures with the target models, thus their traffic are not directly comparable with HaT).
Appendix D Additional Results in the OTSE setting
Addtional results on NinaPro and AD are presented in Figure 9(a) and Figure 9(b) when the target architectures and the portion of labeled data $\gamma$ are varied.
As shown in Figure 9(a) and Figure 9(b), HaT achieves superior or comparable performance in most cases. Although HaT performs slightly worse than MEHLSoup on NinaPro, it significantly outperforms MEHLSoup on the other datasets, likely due to the lower data heterogeneity in NinaPro, which aligns better with MEHLSoup’s approach.